# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

### v2

@@ 4 Benchmark
STHGCN 0.453 0.241 0.197 0.225 0.168 0.223 0.146 0.174 0.250 0.193 0.227 0.250
4.1 POI Recommendation
This benchmark covers the common task of POI recommendation, where the goal is to predict where a user will go next based on their previous check-ins. The input to the model is a trajectory, which is a sequence of places the user has visited. The model is expected to suggest a set of $K$ POIs that the user might visit next. This task can also be framed as re-ranking a set of $N$ candidate POIs in the city based on the user’s recent behavior. This task reflects a typical real-world use case in POI recommendation systems and human mobility modeling. It is a supervised task as the model is trained using all available historical trajectories of the user, allowing it to learn personalized patterns of movement over time. Appendix C provides the details on problem formulation, hyperparameters, experimental setups, and the full results with all evaluation metrics.
Dataset Preparation
By design, our dataset is structured to support consistency and reproducibility, making it straightforward to use for supervised POI recommendation tasks. We adopted the predefined trajectories from the original STD, which has grouped sequences of check-ins into trajectories based on a fixed time interval (see Section 3.1.1). Since all input features in our dataset have been numerically encoded to facilitate model training, we simply used them across all experiments.
For this task, models typically use four kinds of features: (1) social features: user ID; (2) spatial features: POI ID, geographic coordinates; (3) temporal features: check-in timestamp; and (4) categorical features: POI category. Since not all POIs have exact geographic coordinates (see Section 3.1.2), we use the geographic coordinates of their administrative region as a proxy for all POIs.
Models
We evaluated three kinds of models: (1) Markov-based methods, (2) classical deep learning models, and (3) Transformer-based graph neural networks:
  • FPMC [32]: A classical baseline that combines first-order Markov chains with matrix factorization to model personalized next-location predictions.
  • RNN [41], LSTPM [35], and DeepMove [8]: Recurrent neural networks designed to capture sequential dependencies, with varying mechanisms to incorporate spatio-temporal context.
  • GETNext [60] and STHGCN [54]: Transformer-based graph neural networks to model social, spatial, and temporal dependencies.
Evaluation Metrics
We used two popular metrics in POI recommender systems: Acc@k, which checks if the true POI appears in the top-k predicted results, and NDCG@k, which measures the ranking quality of the suggested results.
Results
As shown in Table 3, STHGCN consistently achieves the highest accuracy across all cities, highlighting the effectiveness of Transformer-based GNNs. GETNext often ranks second, performing well in most cities, and older RNN-based models such as LSTPM and DeepMove also remain competitive in a few cities. The best-performing model achieves an average Acc@1 of 22.9%, which is comparable to results from previous studies using datasets of a similar size [7].
We investigated which urban features affect POI recommendation accuracy by computing Spearman correlations between various city features and model performance. We found that category entropy, based on Shannon entropy, strongly correlates with recommendation accuracy ($r=-0.736$), as shown in Fig. 6. Namely, cities with more evenly distributed POI categories and no dominant POI category tend to have lower accuracy and are thus harder to predict. This finding is consistent with the results of prior work on other datasets. Further details are provided in Appendix D.
4.2 Zero-shot POI Recommendation
An increasingly popular approach in POI recommendation leverages the zero-shot capabilities of LLMs, allowing models to perform zero-shot POI recommendation without any additional training. This setting challenges LLMs to generalize to any target city and personalize recommendations for any user. The task mirrors the supervised counterpart: given a user’s trajectory, the model predicts the user’s next destination by ranking a set of $K$ POIs. Appendix E provides the details on problem formulation, prompts, experimental setups, and the full results with all evaluation metrics.
Dataset Preparation
Since LLMs are central for this task, trajectories must be transformed into textual prompts [52, 51]. We adapted the prompts used in [7], which implemented the three LLM methods evaluated in our study. Notably, because LLMs can effectively leverage contextual information, not all input features need to be numerically encoded. The evaluated methods utilize only a subset of features: the timestamp, POI category name, and POI ID.
Methods and Models
We evaluated three LLM-based prompting methods:
  • LLM-Mob [42]: One of the earliest methods to use LLMs for next POI prediction, prompting LLMs with both historical and current (contextual) trajectories.
  • LLM-ZS [1]: A simplified version of LLM-Mob that retains the use of historical and contextual trajectories but simplifies its prompt design.
  • LLM-Move [10]: Extends previous prompting methods by introducing a RAG-like approach, retrieving nearby POIs as candidates, and ranking them by geographic distance to the user’s most recent visit.
To ensure a more robust evaluation, we tested each method using four different LLMs: one closed-source API (Gemini 2.0 Flash [36]) and three open-source models (Qwen 2.5 7B [38], Llama 3.1 8B  [12], and Gemma 2 9B [37]). All open-source models are instruction-tuned variants, quantized to INT4 using AWQ [25], and served through vLLM [21].
Evaluation Metrics
We used the same metrics as in the supervised setting, Acc@k and NDCG@k, for the same reasons: to assess predictive accuracy and to evaluate the quality of the ranking.
Results
As shown in Table 4, LLM-Move [10] performed the best among the three methods, likely due to its well-crafted prompts that supply potential candidate POIs, rather than relying solely on historical or contextual trajectories like LLM-Mob and LLM-ZS. This also supports the use of administrative region coordinates as a proxy when exact POI coordinates are unavailable (e.g., due to privacy issues). Across different LLMs, Gemini 2.0 Flash achieved the highest accuracy across all three prompting strategies, with Qwen 2.5 7B and Gemma 2 9B following closely as strong open-source alternatives. In contrast, we found that Llama 3.1 8B often struggled to follow prompt instructions and frequently produced irrelevant predictions, especially when using the prompts of LLM-Mob and LLM-Move.
Notably, these zero-shot methods outperformed, matched, or came close to the performance of supervised baselines in several cities (e.g., Jakarta, Kuwait City, Moscow), demonstrating their effectiveness even without additional fine-tuning. These results indicate that LLMs can effectively leverage the contextual information included in trajectory prompts to guide their predictions. Since LLMs are pre-trained, they do not require additional training, which reduces both computation time and cost. While inference may require more powerful hardware, it can still be faster overall than training supervised models from scratch.
[…]
Method LLM Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
[…]
Qwen 2.5 7B 0.058 0.035 0.105 0.080 0.030 0.090 0.070 0.030 0.090 0.040 0.035 0.110
Llama 3.1 8B 0.000 0.020 0.055 0.030 0.010 0.030 0.025 0.010 0.030 0.005 0.020 0.005
@@ Results
Qwen 2.5 7B 0.038 0.040 0.065 0.050 0.040 0.080 0.050 0.045 0.095 0.045 0.045 0.120
Llama 3.1 8B 0.077 0.040 0.045 0.060 0.040 0.080 0.055 0.030 0.030 0.060 0.040 0.110
Gemma 2 9B 0.096 0.045 0.105 0.070 0.050 0.080 0.075 0.065 0.075 0.050 0.045 0.110
@@
5 Limitations and Future Work
Firstly, Massive-STEPS is derived from the Semantic Trails dataset and thus inherits its biases and potential errors, which may propagate through downstream tasks. Additionally, the dataset is sparse in several cities, which can impact model training quality and limit cross-city generalization. Secondly, Massive-STEPS focuses solely on trajectories and POI metadata, without including user demographic or social information (e.g., age, social connections). This restricts its applicability for personalized or socially-aware POI recommendation tasks. Thirdly, while our benchmarking covers a wide range of models and cities to emphasize replicability and geographic breadth, we did not perform extensive hyperparameter tuning, which may affect the peak performance of the evaluated models.
While the current version of Massive-STEPS already surpasses existing POI check-in datasets in scale and diversity, we envision expanding it further to include even more cities worldwide. Our dataset creation pipeline is designed to be easily extensible, allowing researchers to integrate additional cities with minimal effort. This allows Massive-STEPS to evolve into a truly global POI check-in dataset.
6 Conclusion
@@
In this paper, we presented the Massive-STEPS dataset to address longstanding limitations in POI recommendation research, particularly the reliance on older, geographically saturated, and non-reproducible check-in datasets. Massive-STEPS offers a large-scale, semantically enriched resource spanning 12 cities across diverse global regions and two time periods, supporting both longitudinal and cross-city analyses. The dataset includes rich semantic information such as venue name, address, category, and coordinates. We also provide benchmark results for both supervised and zero-shot POI recommendation methods, illustrating the dataset’s utility across model types and urban contexts. By releasing Massive-STEPS and our evaluation pipeline publicly, we aim to advance open, reproducible, and globally inclusive research in human mobility and POI recommendation systems. We released our dataset under the Apache 2.0 license.
Acknowledgments and Disclosure of Funding
@@ Appendix A Existing POI Recommendation Datasets
To examine the trend of the usage of POI recommendation datasets, we filtered the comprehensive survey by [68] to extract studies that explicitly mention the cities used in their experiments. The resulting distribution is summarized in Table 5, which shows a strong concentration of studies focused on New York and Tokyo. Additionally, Fig. 1 visualizes the same data, highlighting the uneven distribution of city choices across studies. We also include information on the LBSN platforms used, revealing that Foursquare remains the predominant data source in the field. These findings underscore the need for broader, more inclusive datasets that support evaluation across a wider range of global cities.
@@
Table 5: Overview of POI Recommendation Studies by City and LBSN Platform. This table is adapted from Table IV in the survey by [68] and presents a filtered list of POI recommendation studies that explicitly mention city names and their associated LBSN platforms.
Study Cities LBSN
@@
SSTPMF [5] New York, Tokyo Foursquare, Gowalla
@@
ST-LSTM [69] California, Singapore Brightkite, Foursquare, Gowalla
@@
LSMA [43] New York, San Fransisco, Tokyo Foursquare, Weeplaces
DLAN [46] New York, Tokyo Foursquare
TLR-M [13] New York, Tokyo Foursquare
@@
GETNext [60] New York, Tokyo, California Foursquare, Gowalla
@@
CARAN [17] New York, Tokyo Foursquare, Gowalla
@@
JANICP [70] New York, Tokyo Foursquare, Weeplaces
Li et al. [23] New York, Tokyo Foursquare
@@
AMACF [58] New York, Tokyo Foursquare, Weeplaces
CHA [64] New York, Tokyo Foursquare
HAT [47] Beijing, Shanghai Yelp, others
@@
STAR-HiT [49] New York Foursquare, Gowalla
@@
CAFPR [14] Tokyo, California, Budapest, Melbourne Foursquare
TGAT [19] New York, Tokyo Foursquare
MobGT [50] New York Foursquare, Gowalla
@@
POIBERT [16] Budapest, Delhi, Edinburgh, Glasgow, Osaka, Perth, Toronto Flickr
AutoMTN [30] New York, Tokyo Foursquare
@@
CCDSA [40] New York, Tokyo, San Fransisco Foursquare, Weeplaces
@@
TDGCN [2] Tokyo, California Foursquare, Gowalla, Weeplaces
BayMAN [48] New York Foursquare, Gowalla
@@
ROTAN [11] New York, Tokyo, California Foursquare, Gowalla
@@
STGCN [15] Boston, Chicago, London Gowalla, others
@@
ADQ-GNN [44] New York, Tokyo Foursquare, Gowalla
@@
HS-GAT [67] Boston, Chicago, London Yelp, others
@@
HKGNN [66] New York, Jakarta, Kuala Lumpur, Sao Paulo Foursquare
@@
S2GRec [24] New York, Tokyo Foursquare, Gowalla
@@
GSBPL [39] New York, Tokyo Foursquare, Gowalla
LSPSL [18] New York, Tokyo Foursquare
@@
SCL [3] Florence, Rome, Pisa, Edinburgh, Toronto Flickr
LLM-Move [10] New York, Tokyo Foursquare
@@
LLM4POI [22] New York, Tokyo, California Foursquare, Gowalla
@@
DiffPOI [31] Singapore, New York, Tokyo Foursquare, Gowalla
DSDRec [45] New York, Tokyo Foursquare
@@
Diff-DGMN [72] Istanbul, Jakarta, Sao Paulo, New York, Los Angeles Foursquare

### v3

@@ 4.2.2. Main Results
4.2.3. Longitudinal Experiments and Results
To examine temporal changes in mobility patterns, we split the test sets into two periods (2012-2013 and 2017-2018) and evaluated the same four LLMs with LLM-Move (Feng et al., 2024a), which was the strongest-performing approach in our zero-shot experiments. As shown in Table 6, zero-shot accuracy generally declined in the 2017-2018 period, except for Jakarta and Tangerang, indicating that user trajectories in later years tend to be more challenging to predict. Performance trends varied across cities, highlighting temporal differences in mobility patterns that impact downstream tasks. Across all models, Gemini 2.0 Flash consistently achieved the highest accuracy, demonstrating robust zero-shot capabilities across cities and time.
[…]
Model Bandung Beijing Istanbul Jakarta KC Melbourne Moscow NY Palembang PJ SP Shanghai Sydney Tangerang Tokyo
@@ 4.2.3. Longitudinal Experiments and Results
[…]
XGBoost (BoW) 0.620 0.577 0.600 0.635 0.745 0.695 0.670 0.570 0.625 0.580 0.645 0.600 0.600 0.615 0.565
XGBoost (TF-IDF) 0.660 0.615 0.610 0.640 0.735 0.665 0.650 0.565 0.595 0.580 0.685 0.500 0.585 0.620 0.560
[…]
GPT-4.1 Mini 0.585 0.673 0.615 0.600 0.690 0.585 0.745 0.595 0.605 0.575 0.700 0.565 0.515 0.620 0.550
[…]
4.3. Spatiotemporal Classification and Reasoning
This benchmark assesses whether LLMs can be leveraged for spatiotemporal trajectory classification by providing them with contextual information about a user’s behavior. The task evaluates the model’s ability to capture variations in travel patterns across different cities, given the sequence of POI check-ins as input, and without any additional fine-tuning. Through this setup, we aim to understand how effectively LLMs can reason over spatiotemporal and behavioral cues in user trajectories. Appendix F provides details on problem formulation, prompts, and LLM parameters.
4.3.1. Experimental Setup
This task involves trajectory property classification, specifically predicting whether the final check-in occurs on a weekday or a weekend. Each trajectory is converted into a textual prompt incorporating spatial (city), temporal (check-in time-of-day), and categorical contexts (POI category). Adapting the prompt design from LLM-Mob (Wang et al., 2023c), we instructed the LLM to first reason before making a prediction. This approach allows us to evaluate both classification accuracy and the spatiotemporal reasoning capabilities of LLMs, in line with recent work on spatiotemporal reasoning using LLMs (Quan et al., 2025). Whereas prior approaches rely on models that encode trajectories (Nayak and Pandit, 2023), our method directly leverages the LLM’s ability to process contextual information in natural language. We evaluated four closed-source LLM APIs: Gemini 2.0 Flash (Team and et al., 2024a), GPT-4o Mini, GPT-4.1 Mini, and GPT-5 Nano (OpenAI and et al., 2024a, b), and used Accuracy as our primary metric.
4.3.2. Baseline: Bag-of-POI Categories
As a baseline, we also evaluated classical machine learning classifiers: Logistic Regression, Random Forest, and XGBoost (Chen and Guestrin, 2016). For feature extraction, instead of textual prompting, we propose to simply transform each trajectory into a semantic sequence of POI categories. These sequences are then vectorized using either Bag-of-Words (BoW) or TF-IDF, treating POI categories as ”words” and historical user trajectories as a corpus. Unlike the LLMs, these classical models are trained in a fully supervised manner and are very lightweight.
4.3.3. Main Results
As shown in Table 7, Logistic Regression with TF-IDF achieves the highest mean accuracy (0.658) across the 15 cities, outperforming modern LLMs. The best-performing LLM (Gemini 2 Flash) achieves a mean accuracy of 0.645, performing slightly below classical ML baselines on average. This suggests that simple frequency-based representations of categorical features remain highly competitive for this task. Interestingly, the GPT series generally performs worse, even for models that are more recent than Gemini 2 Flash. For instance, GPT-5 Nano records the lowest mean accuracy, despite being designed to support more advanced reasoning capabilities. Overall, these results highlight current limitations of LLMs in capturing spatiotemporal patterns from trajectory data alone. The stronger performance of classical baselines further indicates that, in a zero-shot setting, LLMs still struggle to reason over spatiotemporal patterns as effectively as traditional statistical methods capture categorical regularities.
5. Conclusion and Limitations
Conclusion
@@
In this paper, we presented Massive-STEPS to address longstanding limitations in POI trajectory modeling, particularly the reliance on older, geographically saturated, and non-reproducible check-in datasets. Massive-STEPS offers a large-scale, semantically enriched POI check-in dataset spanning 15 cities across diverse regions and two time periods, supporting single-city, cross-city, and longitudinal studies. The dataset includes rich semantic information about a POI check-in record, including its name, address, category, and geographic coordinates. We also provide comprehensive benchmarking results for supervised POI recommendation, zero-shot POI recommendation, and spatiotemporal classification, demonstrating the dataset’s versatility across models and tasks. Across these tasks, GNNs performed best for supervised POI recommendation, zero-shot LLMs were competitive in some cities, and classical ML models remained strong for spatiotemporal classification. By releasing our dataset and benchmark pipeline publicly, we aim to advance open, reproducible, and geographically inclusive research in POI modeling and human mobility studies.
Limitations
First, Massive-STEPS is derived from Semantic Trails, thus inheriting its biases, which may propagate through downstream tasks. Additionally, the dataset is sparse in several cities, which can impact model training quality and limit cross-city generalization. Second, Massive-STEPS focuses solely on trajectories and POI metadata, without including user demographic or social information due to privacy considerations. This restricts its applicability for personalized or socially-aware POI recommendation tasks. Third, while our benchmarking covers a wide range of models and cities to emphasize replicability and geographic breadth, we did not perform extensive hyperparameter tuning, which may affect the peak performance of the models. Finally, although Massive-STEPS does not reflect present-day mobility patterns, it was designed to provide an alternative to older datasets and to help bridge the gap toward newer, open, and extensible POI benchmarks.
Acknowledgements.
@@ Appendix A Existing POI Recommendation Datasets
To examine the trend of the usage of POI recommendation datasets, we filtered the comprehensive survey by Zhang et al. (2025) to extract studies that explicitly mention the cities used in their experiments. The resulting distribution is summarized in Table 8, which shows a strong concentration of studies focused on New York and Tokyo. Additionally, Fig. 7 visualizes the same data, highlighting the uneven distribution of city choices across studies. We also include information on the LBSN platforms used, revealing that Foursquare remains the predominant data source in the field. These findings underscore the need for broader, more inclusive datasets that support evaluation across a wider range of global cities.
@@
Table 8. Overview of POI Recommendation Studies by City and LBSN Platform. This table is adapted from Table IV in the survey by Zhang et al. (2025) and presents a filtered list of POI recommendation studies that explicitly mention city names and their associated LBSN platforms.
Study Cities LBSN Dataset Reproducibility Issue
@@
SSTPMF (Davtalab and Alesheikh, 2021) New York, Tokyo Foursquare, Gowalla Gowalla city boundaries not reproducible.
@@
ST-LSTM (Zhao et al., 2018) California, Singapore Brightkite, Foursquare, Gowalla FSQ city boundaries not reproducible. Brightkite and Gowalla not grouped into cities.
@@
LSMA (Wang et al., 2022) New York, San Francisco, Tokyo Foursquare, Weeplaces Weeplaces no longer available.
DLAN (Wu et al., 2024) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
TLR-M (Halder et al., 2021) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
@@
GETNext (Yang et al., 2022c) New York, Tokyo, California Foursquare, Gowalla Only provides preprocessed NYC, missing TKY and CA.
@@
CARAN (Hossain et al., 2022) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
@@
JANICP (Zhong et al., 2022) New York, Tokyo Foursquare, Weeplaces Weeplaces no longer available.
Li et al. (Li et al., 2022a) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
@@
AMACF (Yang et al., 2022a) New York, Tokyo Foursquare, Weeplaces Weeplaces no longer available.
CHA (Zang et al., 2021) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
HAT (Wu et al., 2023) Beijing, Shanghai Yelp, others Used private datasets.
@@
STAR-HiT (Xie and Chen, 2023) New York Foursquare, Gowalla Gowalla not grouped into cities.
@@
CAFPR (Halder et al., 2023) Tokyo, California, Budapest, Melbourne Foursquare Uses POI themepark dataset. POI metadata (lat./lon., category) is missing.
TGAT (Jiang and Wu, 2023) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
MobGT (Xu et al., 2023) New York Foursquare, Gowalla Used private datasets.
@@
POIBERT (Ho and Lim, 2022) Budapest, Delhi, Edinburgh, Glasgow, Osaka, Perth, Toronto Flickr Used private datasets.
AutoMTN (Qin et al., 2022) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
@@
CCDSA (Wang et al., 2023b) New York, Tokyo, San Francisco Foursquare, Weeplaces Weeplaces no longer available.
@@
TDGCN (Cao et al., 2023) Tokyo, California Foursquare, Gowalla, Weeplaces Weeplaces no longer available.
BayMAN (Xia et al., 2023) New York Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
@@
ROTAN (Feng et al., 2024b) New York, Tokyo, California Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
TrajMoE (Han et al., 2025a) Atlanta, Chicago, Seattle, Washington, New York, Los Angeles Unspecified Used private datasets.
UniMove (Han et al., 2025b) Lhasa, Nanchang, Shanghai Unspecified Used private datasets.
@@
STGCN (Han et al., 2020) Boston, Chicago, London Gowalla, others Gowalla city boundaries not reproducible. Used private datasets.
@@
ADQ-GNN (Wang et al., 2021b) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
@@
HS-GAT (Zhang and Ma, 2024) Boston, Chicago, London Yelp, others Yelp not grouped into cities.
@@
HKGNN (Zhang et al., 2024b) New York, Jakarta, Kuala Lumpur, São Paulo Foursquare FSQ city boundaries not reproducible.
@@
S2GRec (Li et al., 2022b) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
@@
GSBPL (Wang et al., 2023a) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
LSPSL (Jiang et al., 2023) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
@@
SCL (Chen and Zhu, 2025) Florence, Rome, Pisa, Edinburgh, Toronto Flickr Preprocessed Flickr dataset no longer available.
LLM-Move (Feng et al., 2024a) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
LLM4POI (Li et al., 2024) New York, Tokyo, California Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
Refine-POI (Li et al., 2025) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
@@
GNPR-SID (Wang et al., 2025) New York, Tokyo, California Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
QT-Mob (Chen et al., 2025) New York, Singapore Foursquare, Private Telco Only FSQ-NYC is publicly available.
@@
DiffPOI (Qin et al., 2023) Singapore, New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
DSDRec (Wang et al., 2024) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
@@
Diff-DGMN (Zuo and Zhang, 2024) Istanbul, Jakarta, São Paulo, New York, Los Angeles Foursquare FSQ city boundaries not reproducible.